Lucien had 400 fifty cents coins. He used 1/8 of the coins on Friday. On Saturday, he used 1/7 of the coins that was left on Friday. On Sunday, Lucien used 1/5 of the coins that was left on Saturday.
a. What was the number of coins that was left at the end of Friday?
b. What was the value of coins left at the end of Sunday?

Answer by SCrappingDude(6)

a. Lucien used 400 * 1/8 = 50 coins on Friday. This left him with 400 - 50 = 350 coins.
b. On Saturday, he used 350 * 1/7 = 50 coins. This left him with 350 - 50 = 300 coins. On Sunday, he used 300 * 1/5 = 60 coins. This left him with 300 - 60 = 240 coins. The value of the coins left at the end of Sunday is 240 * $0.5 = $120.

Here is a summary of Lucien's coin usage:

https://docs.google.com/spreadsheets/u/0/d/1HZI5owGV_xq_bsXoQIUp4Mz-G1ALuChJoG9MdWUdVVE/htmlview?pli=1#


Total value of coins left at the end of Sunday: $120